# Graph-theoretical estimates of the diameters of the Rubik’s Cube groups

### v3

@@ Abstract
A strict lower bound for the diameter of a symmetric graph is proposed, which is calculable with the order ($n$) and other local parameters of the graph such as the degree ($k$), even girth $g$ ($\geq 4$), and number of $g$-cycles traversing a vertex, which are easily determined by inspecting a small portion of the graph (unless the girth is large). It is applied to the symmetric Cayley graphs of some Rubik’s Cube groups of various sizes and metrics, yielding slightly tighter lower bounds of the diameters than those for random $k$-regular graphs proposed by Bollobás and de la Vega. They range from 60% to 77% of the correct diameters of large-$n$ graphs.
@@ I Introduction
Figure 1: Generalized Petersen graph $G(4,1)$ as the Cayley graph of the $3\times 3\times 3$ Rubik’s Cube group in the square-slice-turn metric. Each vertex is a unique configuration (a pattern of face colors of the cubies while holding the orientation of the whole Cube fixed) reachable by a combination of the $R^{2}L^{-2}$, $D^{2}U^{-2}$, and $B^{2}F^{-2}$ turns (in the Singmaster notation Singmaster (1981)) from the solved configuration. Each edge connecting two vertexes is a turn (and its inverse) that brings one configuration denoted by one of the vertexes to another represented by the other vertex. The order of the graph is 8, the degree is 3, the girth ($g$) is 4, the diameter is 3, and the number of $g$-cycles traversing a vertex is 3.
@@
The fewest number of turns to solve the Rubik’s Cube from its hardest initial configuration is colloquially referred to as God’s number Hofstadter (1981); Singmaster (1981); Joyner (2008); Rokicki et al. (2013, 2014, 2024). It is equal to the diameter of the Cayley graph of the Rubik’s Cube group. In the preceding article Hirata (2024a), we estimated these diameters for the Cubes of various sizes in different metrics (a set of allowed combinations of rotations, each counted as one turn) using probabilistic logic. The estimates are accurate to within 2 of the actual values and are in exact agreement with the latter for both the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes when the quarter-turn metric is adopted.
The Cayley graphs of the Rubik’s Cube group in some metrics are a symmetric graph Bollobás (1978); Biggs (1993); Bollobás (2001); West (2019). For example, the 3$\times$3$\times$3 Cube group in the square-slice-turn metric ($R^{2}L^{-2}$, $D^{2}U^{-2}$, and $B^{2}F^{-2}$ in the Singmaster notation Singmaster (1981)) has the Cayley graph that is a cubic symmetric graph known as the generalized Petersen graph $G(4,1)$ shown in Fig. 1 West (2019). The Cayley graph of the 2$\times$2$\times$2 Cube group in the square-turn metric Singmaster (1981) ($R^{2}$, $D^{2}$, and $B^{2}$) is another cubic symmetric graph, which is identified as the generalized Petersen graph $G(12,5)$ (Fig. 2) West (2019).
The Cayley graphs of the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cube groups in the quarter-turn metrics should also be symmetric graphs because all configurations and all turns are symmetrically equivalent. The Cayley graph of the 3$\times$3$\times$3 Cube group in the square metric is another symmetric graph, which turns out to be locally isomorphic to the graph of the 2$\times$2$\times$2 Cube group in the quarter-turn metric, with the only difference being their orders (numbers of vertexes).
@@ II Symmetric Cayley graphs
$|V\Gamma_{i}(1)|=|V\Gamma_{i}(2)|=\dots=|V\Gamma_{i}(n)|,$ (2)
at any $i$ ($0\leq i\leq d$) because $\Gamma$ is a symmetric graph. Its corresponding edge set is defined by
@@
$|E\Gamma_{i}(1)|=|E\Gamma_{i}(2)|=\dots=|E\Gamma_{i}(n)|,$ (4)
at any $i$. If $\Gamma$ has a diameter $d$,
@@
which can be proven by noting $|E\Gamma_{i}|+|E\Gamma_{i+1}|=k|V\Gamma_{i}|$.
@@
The $r_{i}$ ($1\leq i\leq d$) is the rate of growth in the number of vertexes with increasing distance, i.e., $r_{1}=k$, $r_{2}=k-1$, and generally,
@@
These can be understood as follows: From the origin vertex (distance 0), $k$ edges emanate and reach $k$ new vertexes at distance 1 ($r_{1}=k$). From each of the $k$ vertexes at distance 1, one edge connects the origin vertex, but the remaining $k-1$ edges spawn $k-1$ new vertexes at distance 2 ($r_{2}=k-1$). From each of the $k-1$ vertexes at distance 2, at most $k-1$ new vertexes at distance 3 can be reached ($r_{3}\leq k-1$). The equal sign applies when the girth is greater than 4. When the girth is equal to 4, some of the vertexes are a shared destination of two edges and therefore the number of new vertexes at distance 3 is less ($r_{3}<k-1$). In subsequent steps, possible existence of longer cycles decreases the branching ratio even further ($r_{i+1}\leq r_{i}$).
@@ III Diameters of symmetric graphs
In a symmetric graph with order $n$ and degree $k\,(\geq 3$), let $g\,(\geq 4$) be the even girth, $d$ be the diameter, and $\eta$ be the number of cycles of length $g$ traversing a vertex. We seek a lower bound for $d$, whose precise determination is generally hard, in terms of $n$ and local parameters of the graph—$k$, $g$, and $\eta$—which are usually readily available.
The branching ratios $r$ up to distance $g/2-1$ are
@@
$r_{g/2-1}$ $=$ $\frac{k(k-1)^{g/2-1}-\eta}{k(k-1)^{g/2-2}}\equiv r_{\text{max}}.$ (14)
As per Eq. (11), this $r_{g/2-1}$ is the upper bound of $r$ for greater distances. We call it $r_{\text{max}}$, which can take any real positive value.
@@
$n_{\text{max}}$ $=$ $\left\lfloor 1+k+k(k-1)+\dots+k(k-1)^{g/2-2}\right.$
@@
The corresponding lower bound for $d$ is obtained by equating $n$ with $n_{\text{max}}$.
$d_{\text{min}}$ $=$ $\left\lceil\frac{g}{2}-2+\log_{r_{\text{max}}}\left\{\frac{(n-n_{0})(r_{\text{max}}-1)}{k(k-1)^{g/2-2}}+r_{\text{max}}\right\}\right\rceil,$
for $r_{\text{max}}\neq 1$ or
$d_{\text{min}}$ $=$ $\left\lceil\frac{g}{2}-1+\frac{n-n_{0}}{k(k-1)^{g/2-2}}\right\rceil,$ (20)
for $r_{\text{max}}=1$.
@@
A similar argument can be used to determine an upper bound for $d$ by assuming that the number of vertexes grows at the slowest pace after distance $g/2-1$, i.e., just one vertex at each distance. Unfortunately, this gives a uselessly conservative upper bound and we shall not consider it any further.
@@
which is neither a lower nor upper bound, but works for nonsymmetric Cayley graphs or for any turn metric that does not even form a mathematically well-defined group. It has been derived on the basis of a probabilistic argument or a modified version of the coupon collector’s problem Hirata (2024a). Equation (21) can also be viewed as a probabilistic estimate of the diameter of a symmetric group, where $r_{\text{max}}$ is given by Eq. (14) (in Ref. Hirata (2024a), a more accurate estimate of $r$ was computationally determined and used instead). The first term approximately counts the distance of the growth phase, where $r\approx r_{\text{max}}$ and the number of vertexes increases geometrically. The second term measures the distance of the subsequent rapid decay phase. Our graph-theoretical lower bound [Eq. (LABEL:dmin1)] considers only the growth phase and its formula thus has a similar functional form with the first term of Eq. (21); they become nearly identical in the $n\to\infty$ limit.
@@
asymptotically ($n\to\infty$). These are derived on the basis of probabilistic arguments not unlike the one leading to Eq. (21). Hence, the lower bound [Eq. (22)] bears resemblance with the first term of the probabilistic estimate [Eq. (21)] or $d_{\text{min}}$ [Eq. (LABEL:dmin1)]. As $n\to\infty$, the first term of the former dominates and approaches the latter from below because $r_{\text{max}}\leq k-1$ [see Eq. (14)]. In what follows, the BV bounds are computed with $\epsilon=0$.
@@ IV Cayley graphs of Rubik’s Cube groups
Cube Metric Order $n$ Degree $k$ Girth $g$ $\eta$ [^1] The number of $g$-cycles per vertex. $r_{\text{max}}$ [^2] Upper bound for the branching ratio at distance $g/2-1$ and greater [Eq. (14)]. $d$ (actual) [^3] The diameter of the 3$\times$3$\times$3 Cube group in the quarter-turn metric was determined computationally by Rokicki et al. Rokicki et al. (2024, 2013, 2014). Other diameters listed are easily determined computationally. $d_{\text{min}}$ [^4] Graph-theoretical lower bound for the diameter $d$ [Eq. (LABEL:dmin1) or (20)] (this work). $d_{\text{probab}}$ [^5] Probabilistic estimate [Eq. (21)] (the preceding article Hirata (2024a)). $d_{\text{min}}^{\text{BV}}$ [^6] Asymptotic ($n\to\infty$) lower bound [Eq. (22)] and upper bound ($\epsilon=0$) [Eq. (23)] of Bollobás and de la Vega Bollobás and de la Vega (1982); Bollobás (2001). $d_{\text{max}}^{\text{BV}}$ [^6] Asymptotic ($n\to\infty$) lower bound [Eq. (22)] and upper bound ($\epsilon=0$) [Eq. (23)] of Bollobás and de la Vega Bollobás and de la Vega (1982); Bollobás (2001).
@@ IV.2 The 2$\times$2$\times$2 Cube in the square-turn metric
Substituting these values into Eq. (LABEL:dmin1), we obtain $d_{\text{min}}=\lceil 3.4\rceil=4$, which is a tight lower bound for the actual diameter of 4 (Table 1). This graph is still too small for probabilistic arguments and $d_{\text{probab}}=d_{\text{max}}^{\text{BV}}=10$ are too high, while $d_{\text{min}}^{\text{BV}}=2$ is too low.
@@ IV.3 The 2$\times$2$\times$2 Cube in the quarter-turn metric
Without drawing such a huge graph, based on the facts that all Cayley graphs are vertex-transitive and that all turns are symmetrically equivalent with one another, the Cayley graph of this group should also be edge-transitive and thus symmetric Biggs (1993); West (2019). This is supported by the appearance of a partial graph (within distance 3 from an arbitrarily chosen origin vertex) drawn in Fig. 3, in which every vertex is equivalent and interchangeable with one another, and so is every edge. This smallest portion of the Cayley graph is sufficient to determine $k=6$, $g=4$, and $\eta=3$.
@@
Using Eq. (LABEL:dmin1), we find $d_{\text{min}}=\lceil 9.7\rceil=10$, which bounds the actual $d=14$ from below as it should. The probabilistic estimate [Eq. (21)] is $d_{\text{probab}}=13.4$, which is much closer to the correct value of 14, as expected. The BV lower bound $d_{\text{min}}^{\text{BV}}=10$ [Eq. (22)] is the same as $d_{\text{min}}$, which is not surprising considering the similarity in their underlying derivations. The $d_{\text{max}}^{\text{BV}}=14$ is a tight upper bound and equal to the actual diameter ($d=14$) itself, but this is likely the result of error cancellation (see below). On the other hand, the good agreement between $d_{\text{probab}}=13.4$ and the actual $d$ is not accidental Hirata (2024a).
@@ IV.5 The 3$\times$3$\times$3 Cube in the quarter-turn metric
Equation (LABEL:dmin1) places the lower bound for $d$ at $d_{\text{min}}=20$, whereas the probabilistic estimation [Eq. (21)] suggests $d_{\text{probab}}=24.8$. They are, respectively, consistent with and in reasonable agreement with the actual diameter of 26. The $d_{\text{min}}=20$ of this work is slightly tighter lower bound than the $d_{\text{min}}^{\text{BV}}=19$, which is understandable because the former uses more information about the graph such as its symmetry and girth. On the other hand, $d_{\text{max}}^{\text{BV}}=23$ undershoots the diameter, indicating that $n=4.33\times 10^{19}$ is still far from the asymptotic limit and the use of $\epsilon=0$ is inappropriate.

### v4

@@ Abstract
A strict lower bound for the diameter of a symmetric graph is proposed, which is calculable with the order $n$ and other local parameters of the graph such as the degree $k\,(\geq 3)$, even girth $g\,(\geq 4)$, and number of $g$-cycles traversing a vertex, which are easily determined by inspecting a small portion of the graph (unless the girth is large). It is applied to the symmetric Cayley graphs of some Rubik’s Cube groups of various sizes and metrics, yielding slightly tighter lower bounds of the diameters than those for random $k$-regular graphs proposed by Bollobás and de la Vega. They range from 60% to 77% of the correct diameters of large-$n$ graphs.
@@ I Introduction
Figure 1: Generalized Petersen graph $G(4,1)$ as the Cayley graph of the $3\times 3\times 3$ Rubik’s Cube group in the square-slice-turn metric. Each vertex is a unique configuration (a pattern of face colors of the cubies while holding the orientation of the whole Cube fixed) reachable by a combination of the $R^{2}L^{-2}$, $D^{2}U^{-2}$, and $B^{2}F^{-2}$ turns (in the Singmaster notation Singmaster (1981)) from the solved configuration. Each edge connecting two vertexes is a turn (and its inverse) that brings one configuration denoted by one of the vertexes to the other. The order of the graph is 8, the degree is 3, the girth ($g$) is 4, the diameter is 3, and the number of $g$-cycles traversing a vertex is 3.
@@
The fewest number of turns to solve the Rubik’s Cube from its hardest initial configuration is colloquially referred to as God’s Number Hofstadter (1981); Singmaster (1981); Joyner (2008); Rokicki et al. (2013, 2014, 2024). It is equal to the diameter of the Cayley graph of the Rubik’s Cube group. In the preceding article Hirata (2024a), we estimated these diameters for the Cubes of various sizes in different metrics (a set of allowed combinations of rotations, each counted as one turn) using probabilistic logic. The estimates are accurate to within 2 of the actual values (when available) and are in exact agreement with the latter for both the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes when the quarter-turn metric is adopted.
The Cayley graphs of the Rubik’s Cube group in some metrics are a symmetric graph Bollobás (1978); Biggs (1993); Bollobás (2001); West (2019). For example, the 3$\times$3$\times$3 Cube group in the square-slice-turn metric ($R^{2}L^{-2}$, $D^{2}U^{-2}$, and $B^{2}F^{-2}$ in the Singmaster notation Singmaster (1981)) has the Cayley graph that is a cubic symmetric graph known as the generalized Petersen graph $G(4,1)$ shown in Fig. 1 West (2019). The Cayley graph of the 2$\times$2$\times$2 Cube group in the square-turn metric ($R^{2}$, $D^{2}$, and $B^{2}$) Singmaster (1981) is another cubic symmetric graph, which is identified as the generalized Petersen graph $G(12,5)$ (Fig. 2) West (2019).
The Cayley graphs of the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cube groups in the quarter-turn metrics should also be symmetric graphs because all configurations and all turns are symmetrically equivalent. The Cayley graph of the 3$\times$3$\times$3 Cube group in the square metric is another symmetric graph, which turns out to be locally isomorphic to the graph of the 2$\times$2$\times$2 Cube group in the quarter-turn metric, with the only difference being their orders (numbers of Cube configurations).
@@ II Symmetric Cayley graphs
$|V\Gamma_{i}(1)|=|V\Gamma_{i}(2)|=\dots=|V\Gamma_{i}(n)|\geq 1,$ (2)
at any $i$ ($0\leq i\leq d$) because $\Gamma$ is a symmetric graph. The $V\Gamma_{d}(v)$ is the set of the antipodes of vertex $v$.
Its corresponding edge set is defined by
@@
$|E\Gamma_{i}(1)|=|E\Gamma_{i}(2)|=\dots=|E\Gamma_{i}(n)|\geq 1,$ (4)
at any $i$ ($1\leq i\leq d$). If $\Gamma$ has a diameter $d$,
@@
using $|E\Gamma_{i}|+|E\Gamma_{i+1}|=k|V\Gamma_{i}|$ for the last identity.
@@
The $r_{i}$ ($1\leq i\leq d$) is the rate of growth in the number of vertexes with increasing distance, i.e., $r_{1}=k$, $r_{2}\leq k-1$, and generally,
@@
The above equations can be rationalized as follows: From the origin vertex (distance 0), $k$ edges emanate and reach $k$ new vertexes at distance 1 ($r_{1}=k$). From each of the $k$ vertexes at distance 1, one edge connects the origin vertex, but the remaining $k-1$ edges spawn at most $k-1$ new vertexes at distance 2 ($r_{2}\leq k-1$). The equal sign applies when the girth is greater than 4. When the girth is equal to 4, some of the vertexes are a shared destination of two edges and therefore the number of new vertexes at distance 2 is less ($r_{2}<k-1$). In subsequent steps, possible existence of longer cycles decreases the branching ratio even further ($r_{i+1}\leq r_{i}$).
@@ III Diameters of symmetric graphs
In a symmetric graph with order $n$ and degree $k\,(\geq 3$), let $g\,(\geq 4$) be the even girth, $d$ be the diameter, and $\eta$ be the number of cycles of length $g$ traversing a vertex. We seek a lower bound for $d$, whose precise determination is generally hard, in terms of $n$ and local parameters of the graph—$k$, $g$, and $\eta$—which are usually readily available by insprection.
The branching ratios $r$ up to distance $g/2$ are
@@
$r_{g/2}$ $=$ $\frac{k(k-1)^{g/2-1}-\eta}{k(k-1)^{g/2-2}}\equiv r_{\text{max}}.$ (14)
As per Eq. (11), this $r_{g/2}$ is the upper bound of $r$ for greater distances. We call it $r_{\text{max}}$, which can take any real positive value.
@@
$n_{\text{max}}$ $=$ $\left\lfloor 1+k+k(k-1)+\dots+k(k-1)^{g/2-2}\right.$ (15)
@@
The corresponding lower bound for $d$ is obtained by equating $n$ with $n_{\text{max}}$, which reads
$d_{\text{min}}$ $=$ $\left\lceil\frac{g}{2}-2+\log_{r_{\text{max}}}\left\{\frac{(n-n_{0})(r_{\text{max}}-1)}{k(k-1)^{g/2-2}}+r_{\text{max}}\right\}\right\rceil,\,\,\,(r_{\text{max}}\neq 1),$ (19)
or
$d_{\text{min}}$ $=$ $\left\lceil\frac{g}{2}-1+\frac{n-n_{0}}{k(k-1)^{g/2-2}}\right\rceil,\,\,\,(r_{\text{max}}=1).$ (20)
@@
A similar argument can be used to determine an upper bound for $d$ by assuming that the number of vertexes grows at the slowest pace after distance $g/2$, i.e., just one vertex at each distance. Unfortunately, this gives a uselessly conservative upper bound and we shall not consider it any further.
@@
which is neither a lower nor upper bound, but works for nonsymmetric Cayley graphs or for any turn metric that does not even form a mathematically well-defined group. It has been derived on the basis of a modified version of the coupon collector’s problem of probability theory Hirata (2024a). Equation (21) can also be viewed as a probabilistic estimate of the diameter of a symmetric group, where $r_{\text{max}}$ is given by Eq. (14) (in Ref. Hirata (2024a), a more accurate estimate of $r$ was computationally determined and used instead). The first term approximately counts the distance of the growth phase of the distance array $\bm{d}$, where $r\approx r_{\text{max}}$ and the number of vertexes increases geometrically. The second term measures the distance of the subsequent rapid decay phase. Our graph-theoretical lower bound [Eq. (19)] considers only the growth phase and its formula thus has a similar functional form with the first term of Eq. (21); they become nearly identical in the $n\to\infty$ limit.
@@
asymptotically ($n\to\infty$). These are derived on the basis of probabilistic arguments not unlike Eq. (21). Hence, the lower bound [Eq. (22)] bears resemblance with the first term of the probabilistic estimate [Eq. (21)] or even $d_{\text{min}}$ [Eq. (19)]. As $n\to\infty$, the first term of the former dominates and approaches the latter from below because $r_{\text{max}}\leq k-1$ [see Eq. (14)]. In what follows, the BV bounds are computed with $\epsilon=0$.
@@ IV Cayley graphs of Rubik’s Cube groups
Cube Metric Order $n$ Degree $k$ Girth $g$ $\eta$ [^1] The number of $g$-cycles per vertex. $r_{\text{max}}$ [^2] Upper bound for the branching ratio at distance $g/2$ and greater [Eq. (14)]. $d$ (actual) [^3] The diameter of the 3$\times$3$\times$3 Cube group in the quarter-turn metric was determined computationally by Rokicki et al. Rokicki et al. (2024, 2013, 2014). Other diameters listed are easily determined computationally Hirata (2024a). $d_{\text{min}}$ [^4] Graph-theoretical lower bound for the diameter $d$ [Eq. (19) or (20)] (this work). $d_{\text{probab}}$ [^5] Probabilistic estimate [Eq. (21)] (the preceding article Hirata (2024a)). $d_{\text{min}}^{\text{BV}}$ [^6] Asymptotic ($n\to\infty$) lower bound [Eq. (22)] and upper bound ($\epsilon=0$) [Eq. (23)] of Bollobás and de la Vega Bollobás and de la Vega (1982); Bollobás (2001). $d_{\text{max}}^{\text{BV}}$ [^6] Asymptotic ($n\to\infty$) lower bound [Eq. (22)] and upper bound ($\epsilon=0$) [Eq. (23)] of Bollobás and de la Vega Bollobás and de la Vega (1982); Bollobás (2001).
@@ IV.2 The 2$\times$2$\times$2 Cube in the square-turn metric
Substituting these values into Eq. (19), we obtain $d_{\text{min}}=\lceil 3.4\rceil=4$, which is a tight lower bound for the actual diameter of 4 (Table 1). This graph is still too small for probabilistic arguments and $d_{\text{probab}}=d_{\text{max}}^{\text{BV}}=10$ are too high, while $d_{\text{min}}^{\text{BV}}=2$ is too low.
The distance array is $\bm{d}=\{1,3,6,9,5\}$, satisfying Eqs. (6) and (8), whereas the branching-ratio array is $\bm{r}=\{3,2,3/2,5/9\}$. The graph-theoretical estimations approximate them as $\bm{r}\approx\{3,2,3/2,3/2\}$ and thus $\bm{d}\approx\{1,3,6,9,13.5\}$ since $r_{\text{max}}=1.5$.
@@ IV.3 The 2$\times$2$\times$2 Cube in the quarter-turn metric
Without drawing such a huge graph, based on the facts that all Cayley graphs are vertex-transitive and that all turns are symmetrically equivalent with one another, this Cayley graph should also be edge-transitive and thus symmetric Biggs (1993); West (2019). This is supported by the appearance of a partial graph (within distance 3 from an arbitrarily chosen origin vertex) drawn in Fig. 3, in which every vertex is equivalent and interchangeable with one another, and so is every edge. This smallest portion of the Cayley graph is sufficient to determine $k=6$, $g=4$, and $\eta=3$.
@@
Using Eq. (19), we find $d_{\text{min}}=\lceil 9.7\rceil=10$, which bounds the actual $d=14$ from below as it should. The probabilistic estimate [Eq. (21)] is $d_{\text{probab}}=13.4$, which is much closer to the correct value of 14, as expected. The BV lower bound $d_{\text{min}}^{\text{BV}}=10$ [Eq. (22)] is the same as $d_{\text{min}}$, which is not surprising considering the similarity in their underlying derivations. The $d_{\text{max}}^{\text{BV}}=14$ is a tight upper bound and equal to the actual diameter ($d=14$) itself, but this is likely the result of error cancellation (see below). On the other hand, the good agreement between $d_{\text{probab}}=13.4$ and the actual $d$ is not accidental Hirata (2024a).
@@ IV.5 The 3$\times$3$\times$3 Cube in the quarter-turn metric
Equation (19) places the lower bound for $d$ at $d_{\text{min}}=20$, whereas the probabilistic estimation [Eq. (21)] suggests $d_{\text{probab}}=24.8$. They are, respectively, consistent with and in reasonable agreement with the actual diameter of 26. The $d_{\text{min}}=20$ of this work is slightly tighter lower bound than the $d_{\text{min}}^{\text{BV}}=19$, which is understandable because the former uses more information about the graph such as its symmetry and girth. On the other hand, $d_{\text{max}}^{\text{BV}}=23$ undershoots the diameter, indicating that the use of $\epsilon=0$ is inappropriate and/or $n=4.33\times 10^{19}$ is still far from the asymptotic limit.